What is the equation of parabola having its focus at (2,0)and either is end of the latua rectum at(2,2)

Grade:12

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Kindly check and upload correct and complete question better attach image of source of question

Regards
Vikas TU
4 years ago
Dear studentย 
In this parabolaย 
2a= 2
a = 1ย 
Coordinate of vertex = (1,0)
Equation of parabola = (y-0) = 4a(x-x1)
y^2 = 4a(x-1)
a=1
Equation of parabolaย  =y^2 = 4(x-1)
ย